### v1

# Attention Layers Add Into Low-Dimensional Residual Subspaces

Abstract
While transformer models are widely believed to operate in high-dimensional hidden spaces, we show that attention outputs are confined to a surprisingly low-dimensional subspace, where about 60% of the directions account for 99% of the variance–a phenomenon that is induced by the attention output projection matrix and consistently observed across diverse model families and datasets. Critically, we find this low-rank structure as a fundamental cause of the prevalent dead feature problem in sparse dictionary learning, where it creates a mismatch between randomly initialized features and the intrinsic geometry of the activation space. Building on this insight, we propose a subspace-constrained training method for sparse autoencoders (SAEs), initializing feature directions into the active subspace of activations. Our approach reduces dead features from 87% to below 1% in Attention Output SAEs with 1M features, and can further extend to other sparse dictionary learning methods. Our findings provide both new insights into the geometry of attention and practical tools for improving sparse dictionary learning in large language models.
1 Introduction
Over the past years, mechanistic interpretability has shifted from a collection of proof-of-concept tools (Olsson et al., 2022; Wang et al., 2022; Meng et al., 2023; Gould et al., 2023) toward a fast-growing, scale-driven field (Templeton et al., 2024; Ameisen et al., 2025; Lindsey et al., 2025). This transformation is driven by a wave of sparse dictionary learning methods–such as sparse autoencoders (SAEs) and their variants (Cunningham et al., 2023; Bricken et al., 2023b; Lindsey et al., 2024b), transcoders (Dunefsky et al., 2024; Ge et al., 2024), and low-rank sparse attention (He et al., 2025)–that once targeted small models but are now being pushed to larger architectures and wider model families. As these approaches scale in performance and model coverage, they provide increasingly complete and fine-grained explanations of neural network behavior (Lindsey et al., 2024a; Gao et al., 2024).
However, scaling these approaches presents practical difficulties (Templeton et al., 2024; Gao et al., 2024; Mudide et al., 2025). As models and feature dictionaries grow, the number of parameters increases rapidly, driving up computational costs. At the same time, the prevalence of dead features leads to substantial waste in computation and memory (Templeton et al., 2024; Kissane et al., 2024), limiting the efficiency of interpretability methods. In this work, we identify low-rank activation structure as a major driver of dead features (Section 5.1).
In Section 4, we show that attention outputs exhibit a remarkably strong low-rank structure compared to multilayer perceptron (MLP) outputs and residual streams. Through singular value decomposition and intrinsic dimension analyses (Guth et al., 2023; Staats et al., 2025), we demonstrate that this phenomenon holds universally across layers, model families, and datasets, which is consistent with the universality hypothesis (Olah et al., 2020; Chughtai et al., 2023; Gurnee et al., 2024; Wang et al., 2025). We further trace the origin of this low-rank structure to the anisotropy of the output projection matrix $W^{O}$, which compresses the multi-head outputs into a lower-dimensional subspace.
In Section 5, we investigate how the low-rank nature of attention outputs interacts with SAE training. By evaluating the full suite of open-source SAEs from LlamaScope (He et al., 2024), we show that low intrinsic dimension strongly correlates with the number of dead features, suggesting a mismatch between random initialization and the low-dimensional geometry of the activations. To address this, we propose Active Subspace Initialization, which aligns SAE features with the active subspace of activations, substantially reducing dead features while improving reconstruction. Following Lindsey et al. (2024a) and Gao et al. (2024), we conduct scaling experiments, which further reveal that ASI achieves superior reconstruction across feature counts, and when combined with SparseAdam [^1] https://docs.pytorch.org/docs/stable/generated/torch.optim.SparseAdam.html, it achieves the best reconstruction in large scale and reduces dead features from 87% to below 1% in Attention Output SAEs with 1M features trained on Llama-3.1-8B (Dubey et al., 2024).
Furthermore, we show that Active Subspace Init can generalize to sparse replacement models (He et al., 2025; Dunefsky et al., 2024; Ameisen et al., 2025) (Section 5.4). When applied to other sparse dictionary learning methods, our initialization procedure systematically reduces the prevalence of dead parameters across architectures.
2 Related Work
2.1 Low-Rankness in Attention Mechanisms
Prior work has investigated various notions of “low-rankness” within attention mechanisms.: low-rank approximation of attention patterns (Wang et al., 2020; Tay et al., 2020; Raganato et al., 2020), low-rank parameterization for model compression (Noach and Goldberg, 2020; Hu et al., 2022), and the inherent low-rank bottleneck in single-head outputs (Bhojanapalli et al., 2020).
It is important to note that our perspective differs from these prior lines of work. We demonstrate that the multi-head self-attention outputs naturally exhibit a low-rank structure, revealing a distinct and under-explored phenomenon.
2.2 Linear Representation Hypothesis and Sparse Dictionary Learning Methods
The linear representation hypothesis posits that high-level concepts correspond to linear directions in representation space (Arora et al., 2018; Olah et al., 2020; Elhage et al., 2022; Park et al., 2024). Building on this view, a series of sparse dictionary learning methods have been proposed as interpretability tools, including sparse autoencoders and their variants (Cunningham et al., 2023; Bricken et al., 2023b; Lindsey et al., 2024b), transcoders (Dunefsky et al., 2024; Ge et al., 2024), and low-rank sparse attention (He et al., 2025). These approaches aim to decompose activations into combinations of sparsely activated features, while adopting different strategies, depending on their interpretability objectives, to predict or approximate feature activations. Importantly, this hypothesis has been shown to hold across a wide range of model scales (Templeton et al., 2024; Lieberum et al., 2024; He et al., 2024), architectures (Wang et al., 2025), and modalities (Abdulaal et al., 2024).
2.3 Dead Features in Sparse Dictionary Learning Methods
A persistent challenge in sparse dictionary learning methods is the emergence of dead features [^2] Following Bricken et al. (2023b), we define a feature as dead if it never activates over 10 million tokens in this paper. (Templeton et al., 2024; Kissane et al., 2024), which are also referred to as dead units in sparse replacement models (Dunefsky et al., 2024; Ge et al., 2024; He et al., 2025). These features contribute nothing to reconstruction quality, wasting parameters and computation. Existing approaches to mitigate this issue rely on auxiliary loss terms (Gao et al., 2024; Conerly et al., 2025) or resampling strategies (Bricken et al., 2023b) to encourage feature usage.
3 Preliminaries
3.1 Multi-Head Self-Attention and Notations
We consider a Transformer layer with multi-head self-attention (MHSA). Given input representations $X\in\mathbb{R}^{n\times d}$, where $n$ is the number of tokens and $d$ is the hidden size, each attention head $i$ computes:
$Q_{i}=XW^{Q}_{i},\quad K_{i}=XW^{K}_{i},\quad V_{i}=XW^{V}_{i},\quad W^{Q}_{i},W^{K}_{i},W^{V}_{i}\in\mathbb{R}^{d\times d_{h}},$
where $d_{h}=d/H$ is the dimensionality of each head, and $H$ is the total number of heads. The attention weights and head outputs are then given by:
$A_{i}=\mathrm{softmax}\left(\frac{Q_{i}K_{i}^{\top}}{\sqrt{d_{h}}}\right),\quad Z_{i}=A_{i}V_{i}.$
Let $Z=\text{Concat}[Z_{1},\dots,Z_{H}]\in\mathbb{R}^{n\times d}$ denote the concatenated output of all attention heads (Nanda and Bloom, 2022). The final attention output is computed by applying the output projection:
$O=ZW^{O},\quad W^{O}\in\mathbb{R}^{d\times d}.$
This formulation shows that $O$ can be viewed as the sum of low-dimensional outputs from each head, projected into the residual stream space. $O$ serves as the attention block’s contribution to the residual stream.
3.2 TopK Sparse Autoencoders
In this work, we adopt the TopK sparse autoencoder (TopK SAE) variant introduced by Gao et al. (2024). Unlike standard SAEs that impose an $\ell_{1}$ penalty, TopK SAE enforces exact sparsity by keeping only the top-$k$ activations in the latent representation for each input. Formally, given an input vector $x\in\mathbb{R}^{d}$, the encoder produces
$z=\mathrm{TopK}(W_{e}x+b_{e}),$
where $\mathrm{TopK}(v)$ sets to zero all but the largest $k$ entries of $v$. The decoder then reconstructs
$\hat{x}=W_{d}z+b_{d}.$
The model is trained to minimize the reconstruction loss, optionally augmented with an auxiliary penalty to prevent dead latents:
$\mathcal{L}_{\text{TopK-SAE}}=\|x-\hat{x}\|_{2}^{2}+\alpha\cdot\mathcal{L}_{\text{aux}},$
where $\mathcal{L}_{\text{aux}}$ is an optional term designed to penalize latents that never activate over a training period, and $\alpha$ balances reconstruction fidelity and latent utilization.
4 Low-Rank Structure of Attention Outputs
Figure 1: Across layers, model families and datasets, attention outputs exhibit dramatically lower intrinsic dimensions (details in Section 4.1) than residual streams and MLP outputs, showing that the attention layer writing into a low dimensional subspace of residual stream is a universal phenomenon.
We begin by presenting our central empirical finding: in Transformer models, attention outputs consistently display the strongest low-rank structure compared to MLP outputs and residual streams. As shown in Figure 1, attention outputs have a significantly lower intrinsic dimension. This phenomenon is remarkably robust, holding across different intermediate layers, model families and datasets. These observations highlight that the attention block modifies a subspace of the residual stream, while the MLP operates nearly on the full space.
4.1 Quantifying Low-Rankness with Relative Singular Values
We consider activation matrix $\widetilde{A}\in\mathbb{R}^{n\times d}$, where each row corresponds to one token’s activation vector, and $n$ is the number of data points, while $d$ is the dimensionality of the activation space (e.g., the hidden size of the model). Unless otherwise specified, $\widetilde{A}$ refers to the mean-centered activations. We refer readers to Appendix B for more details of activations sources.
To quantify the rank of data, we perform singular value decomposition (SVD) on $\widetilde{A}$:
$\widetilde{A}=U\Sigma V^{\top},$
[…]
To analyze the intrinsic dimension of these activations, we compute the smallest integer $k$ such that:
$\frac{\sum_{i=1}^{k}\sigma_{i}^{2}}{\sum_{i=1}^{r}\sigma_{i}^{2}}\geq\tau,$
for a given threshold $\tau\in(0,1)$. Since SVD yields the optimal low-rank approximation in terms of reconstruction error, this $k$ provides a principled way to assess how concentrated the activations are in a low-dimensional subspace [^3] We use 0.99 as the threshold in main text, results of some other thresholds are provided in Appendix D, with no influence to the conclusion. (Figure 1). We further compute the fraction of delta downstream loss recovered by different number of components. These metrics complement our central findings, offering a numerical characterization of low-rankness.
4.2 Empirical Evidence of Low-Rank Structure
Figure 2: Results for $1M$ samples from SlimPajama (Soboleva et al., 2023) fed into Llama-3.1-8B. (A) The attention outputs are the most low-rank, as indicated by the sharpest decay in singular values. (B) The fraction of loss recovered by different number of components, selected in descending order of singular values.
We draw our findings from three lines of evidence:
Low Intrinsic Dimension of Attention Outputs
Attention outputs exhibit significantly lower intrinsic dimensionality compared to other activation types, with an intrinsic dimension of around 60% of the total dimensionality. In contrast, MLP outputs and the residual streams show much higher intrinsic dimensions above 90% (Figure 1).
Rapid Spectral Decay in Attention Outputs
Attention outputs exhibit a sharply decaying singular value spectrum. This is quantitatively evidenced by the number of components retaining significant energy: only 1174 singular values exceed 5% of the maximum value in attention outputs, compared to 3276 for MLP outputs and 3943 for the residual streams (Figure 2(a)).
Efficient Downstream Loss Recovery
Attention outputs demonstrates superior dimensional efficiency in downstream task performance. Compared to zero ablation, attention outputs requires only 25.0% of dimensions to recover 95% of the downstream loss. This contrasts sharply with MLP outputs and residual streams, which require 78.1% and 85.9% of dimensions respectively to achieve the same recovery level (Figure 2(b)).
More results of these metrics across different layers, models and datasets are shown in Appendix C.
4.3 Low-Rankness of Attention Outputs Results from the Output Projection Matrix
Among all activation types, attention outputs consistently exhibits the most rapid spectral decay. To investigate whether this low-rank structure originates from the attention heads ($Z$), the output projection matrix ($W^{O}$), or their interaction, we perform a decomposition-based analysis.
Recall that the attention output is computed as $O=ZW^{O}$, where $Z\in\mathbb{R}^{n\times d}$ is the concatenated output of all attention heads, and $W^{O}\in\mathbb{R}^{d\times d}$ is a learned linear projection. To understand how the variance in $O$ is shaped, we analyze the variance of $O$ along an arbitrary unit direction $\hat{e}\in\mathbb{R}^{d}$, given by:
[…]
This expression highlights that the variance along $e$ is determined by two factors: the norm of $W^{O}e$ and the variance of $Z$ projected onto the direction $W^{O}\hat{e}$. Specifically, we can rewrite the variance as:
$\mathrm{Var}(O\hat{e})=\mathrm{Var}(Z\hat{v})\cdot\|v\|_{2}^{2}\quad\text{where}\quad v=W^{O}\hat{e},\ \hat{v}=\frac{v}{\|v\|_{2}}.$
We refer to $\mathrm{Var}(Zv)$ as the contribution of Z, capturing how much variance the head output $Z$ provides in that direction, and $\|v\|_{2}^{2}$ as the contribution of $W^{O}$, measuring how much the output projection $W^{O}$ scales or suppresses that direction.
We compute and visualize both quantities across a set of directions aligned with the right singular vectors of attention output, as shown in Figure 3. Our analysis reveals that the low-rank structure of attention outputs $O$ arises primarily from the anisotropy of $W^{O}$, which heavily compresses the output space into a lower-dimensional subspace. From a mechanistic perspective, an intuitive way to see this is that although each attention head contributes a $d_{\text{head}}$-dimensional subspace, the superposition of heads (Jermyn et al., 2024; He et al., 2025) inherently leads to overlaps among these subspaces. We note the output of the $i^{th}$ head as $\text{head}_{i}$. Consequently, the dimension of the MHSA output satisfies
[…]
$\dim\!\left(\bigcup_{i}\text{span}(\text{head}_{i})\right)\;\leq\;\sum_{i}\dim\!\big{(}\text{span}(\text{head}_{i})\big{)}\;=\;d_{\text{head}}\cdot n_{\text{head}}\;(\;=\;d_{\text{model}}\ \text{in standard MHSA)}.$
Figure 4: The number of dead features and the intrinsic dimension of each layer in Llama-3.1-8B, shows a surprising consistency: activations with lower intrinsic dimensions have more dead features.
5 Active Subspace Initialization for Sparse Autoencoders
5.1 Empirical Correlation Between Low-Rank Structure and Dead Features
To explore how low-rankness affects the interpretability of attention, we use the same framework and data as the original study to evaluate the LlamaScope SAEs (He et al., 2024) [^4] Another prominent open-source SAEs, GemmaScope (Lieberum et al., 2024), train their attention SAEs on Z rather than attention output., which provide a complete set of SAEs trained on attention output, MLP output, and residual stream. We find that the number of dead features is strongly related to intrinsic dimensions, as shown in Figure 4. This observation suggests that dead features may stem from a fundamental mismatch between the SAE’s randomly initialized weights and the geometry of the activation space.
5.2 Active Subspace Initialization for Sparse Autoencoders
To investigate this co-occurrence, we train SAEs on the attention output of Llama-3.1-8B at layer 15, initializing the SAE feature directions in the subspace spanned by the top $d_{\text{init}}$ singular vectors of the activations. As shown in Figure 5, we find that constraining the initialization to a lower-dimensional active subspace substantially decreases the number of dead features.
Based on this observation, we propose Active Subspace Initialization (ASI), a lightweight and generalizable strategy for scaling SAEs to high capacities. Let $d$ be the input dimension, $h$ the SAE hidden dimension, and $n$ the number of data points. Given activation matrices $\widetilde{A}\in\mathbb{R}^{n\times d}$, we compute the SVD:
$\widetilde{A}=U\Sigma V^{\top},\quad V\in\mathbb{R}^{d\times d}.$
We select the top $d_{\text{init}}$ right singular vectors to form the active subspace:
$V_{\text{active}}=V_{:,:d_{\text{init}}}\in\mathbb{R}^{d\times d_{\text{init}}}.$
We then initialize the SAE weights directly in this subspace:
$W_{d}\in\text{span}(V_{\text{active}}),\quad W_{e}=W_{d}^{\top}.$
Figure 5: Proportion of dead features and normalized MSE across different subspace dimensions. Random subspaces are used as the baseline, whereas only initialization with the active subspace yields improvement.
Using Active Subspace Initialization offers several benefits:
Figure 6: At a fixed number of features ($n=32768$), Active Subspace Init achieves a better reconstruction-sparsity trade-off than Base TopK and AuxK. A similar trend is observed in its impact on Language Model Downstream Loss. Note: The improvement in downstream loss is less pronounced than that in reconstruction, likely because Active Subspace Init allocates more features to the active subspace, which implicitly enhances reconstruction quality.
Enhanced Sparsity-Reconstruction Frontier Without Additional Compute
It empirically outperforms TopK on the sparsity-reconstruction frontier. It achieves slightly better results compared to the auxiliary loss approach while introducing no additional computational overhead (Figure 6).
Optimal Scaling Characteristics
Our approach demonstrates optimal scaling behavior across various model configurations. On SAEs with an expasion factor of 32 times or more, it can achieve the same or even better performance as the auxiliary loss approach with only half the number of parameters (Section 5.3).
Broad Architectural Applicability
The technique maintains applicability to diverse activation functions and architectural variants, as it operates directly on the intrinsic properties of activations. This generalizability is further explored in Section 5.4.
5.3 Scaling Laws
(a) Loss vs. Number of Parameters.
(b) Number of Dead vs. Number of Parameters.
(c) Loss vs. Number of Alive Parameters.
Figure 7: Scaling results of TopK SAEs and their variants enhanced with AuxK, Active Subspace Init, and SparseAdam–all trained on attention output from Llama-3.1-8B. (A) Loss at convergence across different feature counts: Active Subspace Init consistently achieves lower reconstruction error than TopK and AuxK. Active Subspace Init with SparseAdam achieves the best at large scale. (B) Dead features: Active Subspace Init reduces dead features compared to TopK, but still retains many at extremely large scales. Enhanced with SparseAdam, dead features can be reduced to less than 1%. (C) Loss across different number of alive features: Active Subspace Init achieves the most efficient utilization of alive features, while AuxK shows the lowest efficiency. Details in Section 5.3.
To understand how our method scales, we evaluate performance as the number of SAE features increases from 1616K to 11M, keeping other hyperparameters fixed (details in Appendix E). Figure 7 summarizes the results.
Active Subspace Init improves reconstruction.
As shown in Figure 7(a), the normalized MSE follows a smooth power-law decay with increasing feature count. Across all scales, Active Subspace Init consistently outperforms baseline TopK and AuxK, achieving superior reconstruction at equivalent feature counts.
Caveat: some dead features remain at extremely large scales in Active Subspace Init.
Figure 7(b) shows that, when scaling to extremely large feature counts, Active Subspace Init produces more dead features than AuxK. However, reconstruction performance remains better, indicating that the revived features from AuxK contribute little to actual reconstruction quality(Figure 7(c)).
Use Active Subspace Init with SparseAdam further improves performance.
Prior work (Bricken et al., 2023a) identified stale momentum as a key factor in dead feature formation. Building on this insight, we propose using SparseAdam, an optimizer specifically designed for sparse activation settings. By updating only the moments and parameters corresponding to non-zero gradients, SparseAdam naturally avoids stale momentum and thus mitigates the dead feature issue. As shown in Figures 7(a),  7(b), combining Active Subspace Init with SparseAdam substantially reduces dead features while reaching the lowest reconstruction error. While orthogonal to our initialization method, this choice provides a practical complement that further stabilizes training when scaling SAEs to very large capacities. We discuss more about stale momentum and SparseAdam in Appendix A.
5.4 Generalize to Sparse Replacement Models
Recent work by He et al. (2025) reports that Lorsa, a sparse replacement model for attention layers, exhibits a high proportion of dead parameters. We hypothesize that the low-rank structure of attention outputs contributes significantly to this phenomenon.
To test this, we apply Active Subspace Initialization to Lorsa [^5] Specifically, we initialize the feature directions (corresponding to $W^{O}$ in Lorsa) within the active subspace. Other hyperparameters are the same as He et al. (2025).. This modification reduces the proportion of dead parameters from $68.4\%$ to $40.5\%$ under identical hyperparameter settings (Table 1), while also slightly improving reconstruction error.
Table 1: Effect of Active Subspace Initialization on reducing dead parameters in attention replacement model.
Method Vanilla Active Subspace Init
Dead Parameters (%) 68.4 40.5
Normalized MSE 0.130 0.121
These results indicate that Active Subspace Initialization provides an effective strategy for mitigating dead parameters when training sparse replacement models on low-rank activations. We posit that complete elimination of dead heads may require additional mechanisms, such as the tied initialization used in SAEs to ensure alignment between feature encoding and decoding methods666”Match” means encoder can be initialized to predict relatively accurate feature activation values for decoder.. This approach has been shown to be crucial for reducing dead features in SAEs (Gao et al., 2024), and its absence in Lorsa may limit further improvement.
6 Conclusion
We identified the low-rank structure of attention outputs as a fundamental property of Transformer models and a key cause of dead features in sparse dictionary learning. Our proposed Active Subspace Initialization method addresses this by aligning SAE features with the intrinsic geometry of activations, reducing dead features while improving reconstruction quality. The approach generalizes beyond SAEs to sparse replacement models.
Appendix A Stale Momentum as Another Root Cause of Dead Features
Recent work by Bricken et al. [2023a] identifies stale momentum as a key cause to dead feature formation. Specifically, when a feature remains inactive over training steps, its associated optimizer momentum continues to accumulate. If the feature activates, the stale momentum results in disproportionately large updates, destabilizing training and potentially suppressing that feature permanently.
To directly address this, we adopt SparseAdam, an optimizer tailored for sparse activation settings, designed for more efficient use of compute and memory. SparseAdam updates both parameters and moments only when the corresponding feature is active. This could effectively prevent the harmful accumulation of stale momentum. Empirically, we observe that this change substantially reduces the rate of dead feature formation in large-scale SAE training. We believe that this is a core technique for scaling sparse dictionary methods, as stale momentum is a common problem for them.
Appendix B Activation Sources
The spectral characteristics of activations vary substantially across model architectures, datasets, and positional contexts. Below, we describe the experimental configurations used to support a broad and representative analysis.
Models
We study three large language models of different families–Llama-3.1-8B [^7] https://huggingface.co/meta-llama/Llama-3.1-8B, Qwen3-8B [^8] https://huggingface.co/Qwen/Qwen3-8B, and Gemma-2-9B [^9] https://huggingface.co/google/gemma-2-9b–all based on the Transformer architecture. This allows us to assess the robustness of spectral properties under varying model training configurations.
Datasets
To investigate how dataset diversity affects activation spectra, we select three datasets with varying linguistic and domain characteristics: (1) SlimPajama, an English corpus comprising web text, books, and other sources; (2) RedPajamaGithub, a large-scale code corpus; and (3) CCI3-Data, a Chinese dataset with broad domain coverage.
Activation Positions
Unless otherwise specified, activations are extracted from intermediate layers. For example, in LLaMA-3.1-8B (32 layers), we use activations from layer 15 (zero-indexed). We analyze four types of activations: (1) attention output, (2) MLP output, and (3) residual stream (post layer). For each type, we collect 10 million activation vectors. We empirically verify that this sample size is sufficient to produce stable and reproducible spectral analyses.
Appendix C More Low-Rank Result Across Different Models and Datasets
(a) Activation spectra for many samples from SlimPajama fed into pythia-2.8b.
(b) Activation spectra for many samples from CCI3-Data fed into Qwen3-8B.
(c) Activation spectra for many samples from RedPajamaGithub fed into Qwen3-8B.
Figure 8:
We present relative singular values and fraction of loss recovered for some other model-dataset pairs in Figure 8. Models include pythia-2.8b [^10] EleutherAI/pythia-2.8b. Datasets include RedPajamaGithub [^11] https://huggingface.co/datasets/cerebras/SlimPajama-627B and CCI3-Data [^12] https://huggingface.co/datasets/BAAI/CCI3-Data
(a) Intrinsic dimension with threshold $0.5$.
(b) Intrinsic dimension with threshold $0.9$.
(c) Intrinsic dimension with threshold $0.999$.
(d) Intrinsic dimension with threshold $0.9999$.
Figure 9: Comparison of intrinsic dimensions across different variance thresholds.
Appendix D Different Choose of Variance Threshold for Intrinsic Dimension
We use 0.99 as the variance threshold in the main text. We show other threshold chose make no influence to the conclusion in Figure 9. Attention outputs show low-rank structure consistently.
Appendix E SAE Training Details
E.1 Collecting Activations
We truncate each document to 1024 tokens and prepend a <bos> token to the beginning of each document. During training, we exclude the activations corresponding to the <bos> and <eos> tokens.
It has been observed that activations from different sequence positions within the same document are often highly correlated and may lack diversity. To mitigate this issue, it is common to introduce randomness into the training data. Our shuffling strategy maintains a buffer that is reshuffled whenever the buffer is refilled.
E.2 Initialization and Optimization
The decoder columns $W^{dec}_{:,i}$ are initialized uniformly and normalized to achieve the lowest initial reconstruction loss. We find that the specific initialization norm has little impact, as long as in a reasonable scope. For example, initializing $W^{dec}_{:,i}$ uniformly with a fixed bound, as in Conerly et al. [2025], yields similar results. The encoder weights $W^{enc}$ are initialized as the transpose of $W^{dec}$, while both the encoder bias $b^{enc}$ and decoder bias $b^{dec}$ are set to zero.
This initialization scheme ensures that the SAE begins training with an almost zero reconstruction loss. Such initialization has been widely observed to benefit SAE training.
We train SAEs using the Adam and SparseAdam optimizers, both with $\beta_{1}=0.9$, $\beta_{2}=0.999$, and $\epsilon=10^{-8}$.
E.3 Fixed Hyperparameters in Scaling Law
Model, Dataset, Layer, Pos
Llama-3.1-8B, SlimPajama, 15(index start at 0), attention output.
Sparsity
We empirically set $k=50$ for a reasonable sparsity in scaling laws.
Batch Size
We empirically set $batchsize=4096$, which belows the critical batch size.
Learning Rate
The learning rate for Adam and SparseAdam is sweeped separately in $1\mathrm{e}{-5}$, $2\mathrm{e}{-5}$, $4\mathrm{e}{-5}$, $6\mathrm{e}{-5}$, $8\mathrm{e}{-5}$, $1\mathrm{e}{-4}$, $2\mathrm{e}{-4}$, $4\mathrm{e}{-4}$], and we ultimately use $4\mathrm{e}{-5}$ for Adam and $6\mathrm{e}{-5}$ for SparseAdam.
AuxK
We follow Gao et al. [2024] to set auxiliary loss coefficient $\alpha$ as $\frac{1}{32}$. We sweep the $k_{aux}$ in [256, 512, 1024, 2048] and finally choose 512.
Dimension of Subspace for SAE Initialization
As shown in Figure 5, $d_{init}$ is a hyperparameter with a wide range of sub-optimal value space (from 256 to 2048). We use 768 for all experiments.
Total Tokens
We use $2.5B$ tokens for each SAE training.