### v2

# Attention Layers Add Into Low-Dimensional Residual Subspaces

Abstract
Transformer architectures, and their attention mechanisms in particular, form the foundation of modern large language models. While transformer models are widely believed to operate in high-dimensional hidden spaces, we show that attention outputs are confined to a surprisingly low-dimensional subspace, where about 60% of the directions account for 99% of the variance–a phenomenon that is consistently observed across diverse model families and datasets, and is induced by the attention output projection matrix. Critically, we find this low-rank structure as a key factor of the prevalent dead feature problem in sparse dictionary learning, where it creates a mismatch between randomly initialized features and the intrinsic geometry of the activation space. Building on this insight, we propose a subspace-constrained training method for sparse autoencoders (SAEs), initializing feature directions into the active subspace of activations. Our approach reduces dead features from 87% to below 1% in Attention Output SAEs with 1M features, and can further extend to other sparse dictionary learning methods. Our findings provide both new insights into the geometry of attention and practical tools for improving sparse dictionary learning in large language models.
1 Introduction
Over the past years, mechanistic interpretability has shifted from a collection of proof-of-concept tools (Olsson et al., 2022; Wang et al., 2022; Meng et al., 2023; Gould et al., 2023) toward a fast-growing, scale-driven field (Ameisen et al., 2025; Lindsey et al., 2025). This transformation is driven by a wave of sparse dictionary learning methods–such as sparse autoencoders (SAEs) and their variants (Cunningham et al., 2023; Bricken et al., 2023b; Lindsey et al., 2024b), transcoders (Dunefsky et al., 2024; Ge et al., 2024), and low-rank sparse attention (He et al., 2025)–that once targeted small models but are now being pushed to larger architectures and wider model families (Templeton et al., 2024; Gao et al., 2024; Hazra et al., 2025). As these approaches scale in performance and model coverage, they provide increasingly complete and fine-grained explanations of neural network behavior (Lindsey et al., 2024a; Gao et al., 2024).
However, scaling these approaches presents practical difficulties (Templeton et al., 2024; Gao et al., 2024; Mudide et al., 2025). As models and feature dictionaries grow, the number of parameters increases rapidly, driving up computational costs. At the same time, the prevalence of dead features leads to substantial waste in computation and memory (Templeton et al., 2024; Kissane et al., 2024), limiting the efficiency of interpretability methods. In this work, we identify low-rank activation structure as a major driver of dead features (Section 5.1).
In Section 4, we show that attention outputs exhibit a remarkably strong low-rank structure compared to multilayer perceptron (MLP) outputs and residual streams. Through singular value decomposition and intrinsic dimensionality analyses (Guth et al., 2023; Staats et al., 2025), we demonstrate that this phenomenon holds universally across layers, datasets, and model families—Llama 3.1 (Dubey et al., 2024), Gemma 2 (Rivière et al., 2024), and Qwen 3 (Yang et al., 2025), which is consistent with the universality hypothesis (Olah et al., 2020; Chughtai et al., 2023; Gurnee et al., 2024; Wang et al., 2025). We further trace the origin of this low-rank structure to the anisotropy of the output projection matrix $W^{O}$, which compresses the multi-head outputs into a lower-dimensional subspace.
In Section 5, we investigate how the low-rank nature of attention outputs interacts with SAE training. By evaluating the full suite of open-source SAEs from LlamaScope (He et al., 2024), we show that low intrinsic dimensionality strongly correlates with the number of dead features, suggesting a mismatch between random initialization and the low-dimensional geometry of the activations. Drawing inspiration from Phan et al. (2025)’s principal component initialization for the first network layer, we propose Active Subspace Initialization, which aligns SAE features with the active subspace of activations, substantially reducing dead features while improving reconstruction. Following Lindsey et al. (2024a) and Gao et al. (2024), we conduct scaling experiments, which further reveal that ASI achieves superior reconstruction across feature counts, and when combined with SparseAdam [^1] https://docs.pytorch.org/docs/stable/generated/torch.optim.SparseAdam.html, it achieves the best reconstruction in large scale and reduces dead features from 87% to below 1% in Attention Output SAEs with 1M features trained on Llama-3.1-8B (Dubey et al., 2024).
Furthermore, we show that Active Subspace Init can generalize to sparse replacement models (He et al., 2025; Dunefsky et al., 2024; Ameisen et al., 2025) (Section 5.4). When applied to other sparse dictionary learning methods, our initialization procedure systematically reduces the prevalence of dead parameters across architectures.
2 Related Work
2.1 Low-Rankness in Attention Mechanisms
Prior work has investigated various notions of “low-rankness” within attention mechanisms.: low-rank approximation of attention patterns (Wang et al., 2020; Tay et al., 2020; Raganato et al., 2020), low-rank parameterization for model compression (Noach & Goldberg, 2020; Hu et al., 2022), and the inherent low-rank bottleneck in single-head outputs (Bhojanapalli et al., 2020).
Different from these prior lines of work, we demonstrate that the multi-head self-attention outputs exhibit a low-rank structure, revealing a distinct and under-explored phenomenon.
2.2 Linear Representation Hypothesis and Sparse Dictionary Learning Methods
The linear representation hypothesis suggests that high-level concepts align with linear directions in representation space (Arora et al., 2018; Olah et al., 2020; Elhage et al., 2022; Park et al., 2024). Motivated by this view, various sparse dictionary learning methods have been developed for interpretability, including sparse autoencoders and their variants (Cunningham et al., 2023; Bricken et al., 2023b; Lindsey et al., 2024b), transcoders (Dunefsky et al., 2024; Ge et al., 2024), and low-rank sparse attention (He et al., 2025). These methods decompose activations into sparse feature combinations, while differing in their strategies for predicting or approximating feature activations. The hypothesis has been validated across diverse model scales (Templeton et al., 2024; Lieberum et al., 2024; He et al., 2024), architectures (Wang et al., 2025), and modalities (Abdulaal et al., 2024).
2.3 Dead Features in Sparse Dictionary Learning Methods
A persistent challenge in sparse dictionary learning methods is the emergence of dead features [^2] Following Bricken et al. (2023b), we define a feature as dead if it never activates over 10 million tokens in this paper. (Templeton et al., 2024; Kissane et al., 2024), which are also referred to as dead units in sparse replacement models (Dunefsky et al., 2024; Ge et al., 2024; He et al., 2025). These features contribute nothing to reconstruction quality, wasting parameters and computation. Existing approaches to mitigate this issue rely on auxiliary loss terms (Gao et al., 2024; Conerly et al., 2025) or resampling strategies (Bricken et al., 2023b) to encourage feature usage.
2.4 PCA-Inspired Network Initialization
A common practice applies PCA to input data for dimensionality reduction before network training (Hastie et al., 2009; Montavon et al., 2012; Jolliffe, 1986; Bishop & Nasrabadi, 2007). Recently, Phan et al. (2025) proposed PCsInit, which initializes the first layer weights of networks with top principal components of data—embedding the PCA transform directly into the network. This provides the model with a superior parameter set (Gu et al., 2025), boosting performance by construction.
3 Preliminaries
3.1 Multi-Head Self-Attention and Notations
We consider a Transformer block with multi-head self-attention (MHSA) (Vaswani et al., 2017). Given input activations $X\in\mathbb{R}^{n\times d}$, where $n$ is the token count and $d$ is the model hidden size, each attention head $i$ computes:
$Q_{i}=XW^{Q}_{i},\quad K_{i}=XW^{K}_{i},\quad V_{i}=XW^{V}_{i},\quad W^{Q}_{i},W^{K}_{i},W^{V}_{i}\in\mathbb{R}^{d\times d_{h}},$
where $d_{h}=d/H$ is the dimensionality of each head, and $H$ is the total number of heads. The attention weights and head outputs are then given by:
$A_{i}=\mathrm{softmax}\left(\frac{Q_{i}K_{i}^{\top}}{\sqrt{d_{h}}}\right),\quad Z_{i}=A_{i}V_{i}.$
Let $Z=\text{Concat}[Z_{1},\dots,Z_{H}]\in\mathbb{R}^{n\times d}$ denote the concatenated output of all attention heads (Nanda & Bloom, 2022). The final attention output is computed by applying the output projection:
$O=ZW^{O},\quad W^{O}\in\mathbb{R}^{d\times d}.$
This formulation shows that $O$ can be viewed as the sum of low-dimensional outputs from each head, projected into the residual stream space. $O$ is the attention block’s contribution to the residual stream.
3.2 TopK Sparse Autoencoders
In this work, we adopt the TopK sparse autoencoder (TopK SAE) introduced by Gao et al. (2024). Unlike standard SAEs that impose an $\ell_{1}$ penalty, TopK SAE enforces exact sparsity by keeping only the top-$k$ activations in the latent representation for each input. Formally, given an input vector $x\in\mathbb{R}^{d}$, the encoder produces
$z=\mathrm{TopK}(W_{e}x+b_{e}),$
where $\mathrm{TopK}(v)$ sets to zero all but the largest $k$ entries of $v$. The decoder then reconstructs
$\hat{x}=W_{d}z+b_{d}.$
The model is trained to minimize the reconstruction loss, optionally augmented with an auxiliary penalty to prevent dead latents:
$\mathcal{L}_{\text{TopK-SAE}}=\|x-\hat{x}\|_{2}^{2}+\alpha\cdot\mathcal{L}_{\text{aux}},$
where $\mathcal{L}_{\text{aux}}$ is an optional term designed to penalize latents that never activate over a training period, and $\alpha$ balances reconstruction fidelity and latent utilization.
4 Low-Rank Structure of Attention Outputs
Figure 1: Across layers, model families and datasets, attention outputs exhibit dramatically lower intrinsic dimensions (details in Section 4.1) than residual streams and MLP outputs, showing that the attention layer writing into a low dimensional subspace of residual stream is a universal phenomenon.
We begin by presenting our central empirical finding: in Transformer models, attention outputs consistently display the strongest low-rank structure compared to MLP outputs and residual streams. As shown in Figure 1, attention outputs have a significantly lower intrinsic dimension. This phenomenon is remarkably robust, holding across different intermediate layers, model families and datasets. These observations highlight that the attention block modifies a subspace of the residual stream, while the MLP operates nearly on the full space.
4.1 Quantifying Low-Rankness with Relative Singular Values
We consider activation matrix $\widetilde{A}\in\mathbb{R}^{n\times d}$, where each row corresponds to one token’s activation vector, and $n$ is the number of data points, while $d$ is the dimensionality of the activation space (e.g., the hidden size of the model). Unless otherwise specified, $\widetilde{A}$ refers to the mean-centered activations. We refer readers to Appendix D for more details of activations sources.
To quantify the rank of data, we perform singular value decomposition (SVD) on $\widetilde{A}$:
$\widetilde{A}=U\Sigma V^{\top},$
[…]
To analyze the intrinsic dimension of these activations, we compute the smallest integer $k$ such that:
$\frac{\sum_{i=1}^{k}\sigma_{i}^{2}}{\sum_{i=1}^{r}\sigma_{i}^{2}}\geq\tau,$
for a given threshold $\tau\in(0,1)$. Since SVD yields the optimal low-rank approximation in terms of reconstruction error, this $k$ provides a principled way to assess how concentrated the activations are in a low-dimensional subspace [^3] We use 0.99 as the threshold in main text, results of some other thresholds are provided in Appendix F, with no influence to the conclusion. (Figure 1). We further compute the fraction of delta downstream loss recovered by different number of components (Section 4.2). These metrics complement our central findings, offering a numerical characterization of low-rankness.
For each type of activations, we collect 10 million activation vectors [^4] In rare cases, outlier activations inflate variance along some directions, making variance-based dimensionality estimates unreliable; to address this, we filter activations with a norm exceeding $5\sigma$ from the mean.. We empirically verify that this sample size is sufficient to produce stable and reproducible singular spectrum analysis.
4.2 Empirical Evidence of Low-Rank Structure
Figure 2: Results for 1010M tokens from SlimPajama (Soboleva et al., 2023) fed into Llama-3.1-8B. (a) The attention outputs are the most low-rank, as indicated by the sharpest decay in singular values. (b) Fraction of loss recovered using varying numbers of top singular components.
We draw our findings from three lines of evidence:
Low Intrinsic Dimension of Attention Outputs
Attention outputs has an intrinsic dimension of around 60% of the total dimensionality. In contrast, MLP outputs and the residual streams show much higher intrinsic dimensions above 90% (Figure 1).
Rapid Singular Spectral Decay in Attention Outputs
This is quantitatively evidenced by the number of components retaining significant energy: only 74.7% singular values exceed 1% of the maximum in attention outputs, versus 100.0% for MLP outputs and residual streams (Figure 2(a)).
Efficient Downstream Loss Recovery
Compared to zero ablation, attention outputs requires only 39.1% of dimensions to recover over 99% of the downstream loss, versus 95.3% and 96.9% for MLP outputs and residual streams (Figure 2(b)).
More results of these metrics across different layers, models and datasets are shown in Appendix E.
4.3 Low-Rankness of Attention Outputs Results from the Output Projection Matrix
Among all activation types, attention outputs consistently exhibits the most rapid spectral decay. To investigate whether this low-rank structure originates from the attention heads ($Z$), the output projection matrix ($W^{O}$), or their interaction, we perform a decomposition-based analysis.
Recall that the attention output is computed as $O=ZW^{O}$, where $Z\in\mathbb{R}^{n\times d}$ is the concatenated output of all attention heads, and $W^{O}\in\mathbb{R}^{d\times d}$ is a learned linear projection. To understand how the variance in $O$ (singular value spectra of O) is shaped, we analyze the variance of $O$ along an arbitrary unit direction $\hat{e}\in\mathbb{R}^{d}$, given by:
[…]
This expression highlights that the variance along $e$ is determined by two factors: the norm of $W^{O}e$ and the variance of $Z$ projected onto the direction $W^{O}\hat{e}$. Specifically, we can rewrite the variance as:
$\mathrm{Var}(O\hat{e})=\mathrm{Var}(Z\hat{v})\cdot\|v\|_{2}^{2}\quad\text{where}\quad v=W^{O}\hat{e},\ \hat{v}=\frac{v}{\|v\|_{2}}.$
We refer to $\mathrm{Var}(Zv)$ as the contribution of Z, capturing how much variance the head output $Z$ provides in that direction, and $\|v\|_{2}^{2}$ as the contribution of $W^{O}$, measuring how much the output projection $W^{O}$ scales or suppresses that direction.
We compute and visualize both quantities across a set of directions aligned with the right singular vectors of attention output, as shown in Figure 3. Our analysis reveals that the low-rank structure of attention outputs $O$ arises primarily from the anisotropy of $W^{O}$, which heavily compresses the output space into a lower-dimensional subspace. From a mechanistic perspective, an intuitive way to see this is that although each attention head contributes a $d_{\text{head}}$-dimensional subspace, the superposition of heads (Jermyn et al., 2024; He et al., 2025) inherently leads to overlaps among these subspaces. We note the output of the $i^{th}$ head as $\text{head}_{i}$. Consequently, the dimension of the MHSA output satisfies
[…]
$\dim\!\left(\bigcup_{i}\text{span}(\text{head}_{i})\right)\;\leq\;\sum_{i}\dim\!\big(\text{span}(\text{head}_{i})\big)\;=\;d_{\text{head}}\cdot n_{\text{head}}\;(\;=\;d_{\text{model}}\ \text{in standard MHSA)}.$
5 Active Subspace Initialization for Sparse Autoencoders
5.1 Empirical Correlation Between Low-Rank Structure and Dead Features
Figure 4: The number of dead features (left) and the intrinsic dimension (right) of each activation in Llama-3.1-8B, shows a surprising consistency: activations with lower intrinsic dimensions have more dead features, corresponding to all layers of attention output and last two layers of MLP output.
To explore how low-rankness affects the interpretability of attention, we use the same framework and data as the original study to evaluate the LlamaScope SAEs (He et al., 2024), which provide a complete set of SAEs trained on attention output, MLP output, and residual stream  [^5] Another prominent open-source SAEs, GemmaScope (Lieberum et al., 2024), train their attention SAEs on Z rather than attention output.. We find that the number of dead features is strongly related to intrinsic dimensions, as shown in Figure 4. This observation suggests that dead features may stem from the low-rank geometry of the activation space.
5.2 Active Subspace Initialization for Sparse Autoencoders
Based on this observation, we propose Active Subspace Initialization (ASI), a lightweight and generalizable strategy for scaling SAEs to high capacities. Let $d$ denote the input dimension, $h$ the hidden dimension of the SAE, and $n$ the number of data points. Given activation matrices $\widetilde{A}\in\mathbb{R}^{n\times d}$, we compute the singular value decomposition (SVD):
$\widetilde{A}=U\Sigma V^{\top},\quad V\in\mathbb{R}^{d\times d},$
where $U\in\mathbb{R}^{n\times d}$ is the left singular vector matrix, $\Sigma\in\mathbb{R}^{d\times d}$ is the diagonal matrix of singular values, and $V\in\mathbb{R}^{d\times d}$ contains the right singular vectors.
We select the top $d_{\text{init}}$ right singular vectors to form the active subspace and then initialize the SAE weights directly in this subspace:
$V_{\text{active}}=V_{:,:d_{\text{init}}}\in\mathbb{R}^{d\times d_{\text{init}}},\quad W_{D}\in\text{span}(V_{\text{active}}),\quad W_{E}=W_{D}^{\top},$
where $W_{E}$ is the encoder weight matrix and $W_{D}$ is the decoder weight matrix. Intuitively, ASI aligns the initial SAE parameters with the active directions of the data, ensuring that SAEs start in a meaningful low-dimensional subspace. As $d_{\text{init}}$ decreases from the full space dimension [^6] Setting $D_{init}$ equal to the full space dimension is equivalent to not using Active Subspace Initialization. within a certain range, the number of dead features in the SAE rapidly drops, with a corresponding improvement in Delta LM loss [^7] This metric is defined as the difference between the original language model loss and the loss when the SAE is inserted at the corresponding position, evaluated over 1 million tokens. (Figure 5). Further ablation studies and implementation details are contained in Appendix B and Appendix I, respectively.
Figure 5: After using ASI, proportion of dead features (left), normalized MSE (mid) and Delta LM loss (right) across different subspace dimensions for activations with a full space dimension of 4096.
Figure 6: At a fixed number of features ($n=32768$), Active Subspace Init achieves a better reconstruction-sparsity trade-off than TopK and AuxK. A similar trend is observed in its impact on Delta LM Loss. Note: The improvement in downstream loss is less pronounced than that in reconstruction, likely because Active Subspace Init allocates more features to the active subspace, which implicitly enhances reconstruction quality.
Using Active Subspace Initialization offers several benefits:
Reduced dead features and Enhanced Sparsity-Reconstruction Frontier Without Additional Compute
It achieves near-zero dead features and slightly superior results compared to the auxiliary loss approach (AuxK), at no additional computational cost of the same order. (Figure 6).
Optimal Scaling Characteristics
Our approach demonstrates optimal scaling behavior across various SAE training methods. It outperforms TopK and AuxK in any evaluated scale, from 1616K to 11M features (Section 5.3).
General Applicability
The technique maintains applicability to diverse architectural variants and activation functions, as it operates directly on the intrinsic properties of activations. This generalizability is further explored in Section 5.4 and Appendix H.
5.3 Scaling Laws
To understand how our method scales, we evaluate performance as the number of SAE features increases from 1616K to 11M, keeping other hyperparameters fixed (details in Appendix G.5).
(a) Loss vs. Number of Parameters.
(b) Number of Dead vs. Number of Parameters.
(c) Loss vs. Number of Alive Parameters.
Figure 7: Scaling results of TopK SAEs and their variants enhanced with AuxK, Active Subspace Init, and SparseAdam–all trained on attention output from Llama-3.1-8B. (A) Loss at convergence across different feature counts: Active Subspace Init consistently achieves lower reconstruction error than TopK and AuxK. Active Subspace Init with SparseAdam achieves the best at large scale. (B) Dead features: Active Subspace Init reduces dead features compared to TopK, but still retains many at extremely large scales. Enhanced with SparseAdam, dead features can be reduced to less than 1%. (C) Loss across different number of alive features: Active Subspace Init achieves the most efficient utilization of alive features, while AuxK shows the lowest efficiency. Details in Section 5.3.
Active Subspace Init improves reconstruction.
As shown in Figure 7(a), Active Subspace Init consistently outperforms TopK and AuxK across all scales.
Caveat: some dead features remain at extremely large scales in Active Subspace Init.
Figure 7(b) shows that, when scaling to extremely large feature counts, Active Subspace Init produces more dead features than AuxK. However, reconstruction performance remains better, indicating that the revived features from AuxK contribute little to actual reconstruction quality(Figure 7(c)).
Use Active Subspace Init with SparseAdam further improves performance.
Prior work (Bricken et al., 2023a) identified stale momentum as a key factor in dead feature formation. Building on this insight, we propose using SparseAdam, an optimizer specifically designed for sparse activation settings. By updating only the moments and parameters corresponding to non-zero gradients, SparseAdam naturally avoids stale momentum and thus mitigates the dead feature issue. As shown in Figures 7(a),  7(b), combining Active Subspace Init with SparseAdam substantially reduces dead features while reaching the lowest reconstruction error. While orthogonal to our initialization method, this choice provides a practical complement that further stabilizes training when scaling SAEs to very large capacities. We discuss more about stale momentum and SparseAdam in Appendix C.
5.4 Generalize to Sparse Replacement Models
Recent work by He et al. (2025) reports that Lorsa, a sparse replacement model for attention layers, exhibits a high proportion of dead parameters. We hypothesize that the low-rank structure of attention outputs contributes significantly to this phenomenon.
To evaluate this, we apply Active Subspace Initialization to Lorsa. This modification reduces the proportion of dead parameters significantly under identical hyperparameter settings (Figure 8), while also improving the reconstruction. Specifically, we initialize the decoder (corresponding to $W^{O}$ in Lorsa) within the active subspace of each original MHSA head, while the encoder (corresponding to $W^{V}$ in Lorsa) is aligned with the corresponding input-side directions of the decoder. Further implementation details are provided in Appendix I.
Figure 8: Effect of Active Subspace Initialization on reducing dead parameters in attention replacement model.
6 Discussion
When to Use Active Subspace Initialization
For activation sites that do not exhibit clear low-rank structures (e.g., residual stream), our method yields only limited gains (Appendix B.2). The singular value spectrum provides a more reliable indicator: when the spectrum decays rapidly in its tail, applying ASI is likely to be effective.
Causality between Low-Rank Structure and Dead Features
We observed a strong correlation between low-rank structure and the emergence of dead features (Section 5), but the causal mechanism remains unclear. It may be tied to optimization dynamics or feature competition, yet a precise explanation is lacking. A deeper understanding of this relationship could also explain why ASI is less effective under high-$\ell_{0}$ warm-up regimes (Appendix H). We leave this important line of inquiry to future work.
7 Conclusion
We identified the low-rank structure of attention outputs as a fundamental property of Transformer models and a key cause of dead features in sparse dictionary learning. Our proposed Active Subspace Initialization method addresses this by aligning SAE features with the intrinsic geometry of activations, reducing dead features while improving reconstruction quality. The approach generalizes beyond SAEs to sparse replacement models.
Appendix A Declaration of LLM Usage
We acknowledge the use of large language models (LLMs) to assist in the preparation of this manuscript. Specifically, ChatGPT and Claude were utilized during the writing and coding process for the exclusive purpose of improving text clarity, grammar, overall fluency, and for generating boilerplate code structures. All ideas, theoretical developments, experimental designs, results, analyses, and scientific conclusions are entirely our own. The LLMs acted solely as assistive tools and were not involved in any aspect of the intellectual or scientific work.
Appendix B Ablation Study
B.1 Active Subspace Init vs Random Subspace Init
We employ random subspace initialization as a baseline and observe that it consistently degrades SAE training across all metrics, as shown in Figure 9.
Figure 9: For activations with a full space dimension of 4096, proportion of dead features (left), normalized MSE (mid) and Delta LM loss (right) across different subspace dimensions. Random subspaces are used as the baseline, whereas only initialization with the active subspace yields improvement.
B.2 Apply on Near-Full-Rank Activation
We also apply Active Subspace Initialization (ASI) to near-full-rank activations, such as those in the residual stream, to evaluate its generality. When training an SAE on the post-layer-15 residual stream of Llama-3.1-8B, we find ASI yields minimal gains (Figure 10). This is consistent with our expectation, as these activations inherently exhibit a lower rate of dead features even with standard initialization.
Figure 10: For activations with a full space dimension of 4096, proportion of dead features (left), normalized MSE (mid) and Delta LM loss (right) across different subspace dimensions. Random subspaces are used as the baseline, whereas only initialization with the active subspace yields improvement.
B.3 Choice of the Initial Dictionary Size $D_{init}$
As shown in Figure 5, $D_{init}$ is a hyperparameter with a wide range of acceptable values (from 256 to 2048). We hypothesize that the performance degradation at very low values occurs because an excessively small dictionary fails to span the subspace containing critical information needed for effective reconstruction.
Appendix C Stale Momentum as Another Root Cause of Dead Features
Recent work by Bricken et al. (2023a) identifies stale momentum as a key cause to dead feature formation. Specifically, when a feature remains inactive over training steps, its associated optimizer momentum continues to accumulate. If the feature activates, the stale momentum results in disproportionately large updates, destabilizing training and potentially suppressing that feature permanently.
To directly address this, we adopt SparseAdam, an optimizer tailored for sparse activation settings, designed for more efficient use of compute and memory. SparseAdam updates both parameters and moments only when the corresponding feature is active. This could effectively prevent the harmful accumulation of stale momentum. Empirically, we observe that this change substantially reduces the rate of dead feature formation in large-scale SAE training. We believe that this is a core technique for scaling sparse dictionary methods, as stale momentum is a common problem for them.
Appendix D Activation Sources
The spectral characteristics of activations vary substantially across model architectures, datasets, and positional contexts. Below, we describe the experimental configurations used to support a broad and representative analysis.
Models
We study three large language models of different families–Llama-3.1-8B [^8] https://huggingface.co/meta-llama/Llama-3.1-8B, Qwen3-8B [^9] https://huggingface.co/Qwen/Qwen3-8B, and Gemma-2-9B [^10] https://huggingface.co/google/gemma-2-9b–all based on the Transformer architecture. This allows us to assess the robustness of spectral properties under varying model training configurations.
Datasets
To investigate how dataset diversity affects activation spectra, we select three datasets with varying linguistic and domain characteristics: (1) SlimPajama, an English corpus comprising web text, books, and other sources; (2) RedPajamaGithub, a large-scale code corpus; and (3) CCI3-Data, a Chinese dataset with broad domain coverage.
Activation Positions
Unless otherwise specified, activations are extracted from model intermediate layers. For example, in LLaMA-3.1-8B (32 layers), we use activations from layer 15 (zero-indexed). We analyze three types of activations: (1) attention output, (2) MLP output, and (3) residual stream (post layer).
Appendix E More Low-Rank Result Across Different Models and Datasets
(a) Activation spectra for many samples from SlimPajama fed into pythia-2.8b.
(b) Activation spectra for many samples from CCI3-Data fed into Qwen3-8B.
(c) Activation spectra for many samples from RedPajamaGithub fed into Qwen3-8B.
Figure 11:
We present relative singular values and fraction of loss recovered for some other model-dataset pairs in Figure 11. Models include pythia-2.8b [^11] https://huggingface.co/EleutherAI/pythia-2.8b. Datasets include RedPajamaGithub [^12] https://huggingface.co/datasets/cerebras/SlimPajama-627B and CCI3-Data [^13] https://huggingface.co/datasets/BAAI/CCI3-Data
(a) Intrinsic dimension with threshold $0.5$.
(b) Intrinsic dimension with threshold $0.9$.
(c) Intrinsic dimension with threshold $0.999$.
(d) Intrinsic dimension with threshold $0.9999$.
Figure 12: Comparison of intrinsic dimensions across different variance thresholds.
Appendix F Different Choose of Variance Threshold for Intrinsic Dimension
We use 0.99 as the variance threshold in the main text. We show other threshold chose make no influence to the conclusion in Figure 12. Attention outputs show low-rank structure consistently.
Appendix G SAE Training Details
G.1 Collecting Activations
We truncate each document to 1024 tokens and prepend a <bos> token to the beginning of each document. During training, we exclude the activations corresponding to the <bos> and <eos> tokens.
It has been observed that activations from different sequence positions within the same document are often highly correlated and may lack diversity. To mitigate this issue, it is common to introduce randomness into the training data. Our shuffling strategy maintains a buffer that is reshuffled whenever the buffer is refilled.
G.2 Initialization
The decoder columns $W^{dec}_{:,i}$ are initialized uniformly, and the optimal norm for them is found through a grid search to minimize the initial reconstruction loss. We find that the specific initialization norm has little impact, as long as in a reasonable scope. For example, initializing $W^{dec}_{:,i}$ uniformly with a fixed bound, as in Conerly et al. (2025), yields similar results. The encoder weights $W^{enc}$ are initialized as the transpose of $W^{dec}$, while both the encoder bias $b^{enc}$ and decoder bias $b^{dec}$ are set to zero.
G.3 Optimization
We train SAEs using the Adam optimizers, both with $lr=2*10^{-4}$, $batchsize=32768$, $\beta_{1}=0.9$, $\beta_{2}=0.999$, and $\epsilon=10^{-8}$. The differences in scaling law experiments are described in the Appendix G.5.
G.4 Jumprelu SAEs
We trained JumpReLU SAEs (Rajamanoharan et al., 2024) under two distinct hyperparameter configurations: one maintaining a consistently low $\ell_{0}$ value throughout training, and another where $\ell_{0}$ is gradually decreased from a higher initial value. Unless otherwise specified, all JumpReLU SAEs were trained using the same settings as Conerly et al. (2025), which corresponds to the latter configuration. The key modifications for the former setting are as follows: (1) we initialized the encoder bias to zero instead of applying the heuristic that equalizes feature activation counts at initialization, and (2) we kept the sparsity coefficient fixed rather than employing a global warm-up schedule. As a result, the $\ell_{0}$ sparsity level started at a relatively low value early in training. This design is critical to our approach: we observed that if the model remains in a high-$\ell_{0}$ regime (e.g., on the order of $d_{\text{model}}/2$) for an extended period before sparsity increases, the feature directions tend to drift away from the active subspace during this phase, thereby diminishing the effectiveness of our method (Appendix H.1).
G.5 Fixed Hyperparameters in Scaling Law
Model, Dataset, Layer, Pos
Llama-3.1-8B, SlimPajama, 15(index start at 0), attention output.
Sparsity
We empirically set $k=50$ for a reasonable sparsity in scaling laws.
Batch Size
We empirically set the batch size to 40964096—below the critical batch size—which differs from the setting of $32768$ used in other experiments in this paper. This reduction is necessary because training SAEs with 1 million features requires substantial memory, posing significant computational challenges.
Learning Rate
The learning rate for Adam and SparseAdam is sweeped separately in $1\mathrm{e}{-5}$, $2\mathrm{e}{-5}$, $4\mathrm{e}{-5}$, $6\mathrm{e}{-5}$, $8\mathrm{e}{-5}$, $1\mathrm{e}{-4}$, $2\mathrm{e}{-4}$, $4\mathrm{e}{-4}$], and we ultimately use $4\mathrm{e}{-5}$ for Adam and $6\mathrm{e}{-5}$ for SparseAdam. We employ a three-phase learning rate schedule consisting of a linear warm-up, a stable phase, and a linear decay. The learning rate increases linearly from zero to its maximum value over the first 500 steps, remains constant during the intermediate phase, and then decays linearly to 1% of the maximum value over the final 20% of the total training steps.
AuxK
We follow Gao et al. (2024) to set auxiliary loss coefficient $\alpha$ as $\frac{1}{32}$. We sweep the $k_{aux}$ in [256, 512, 1024, 2048] and finally choose 512.
Dimension of Subspace for SAE Initialization
We use 768 for all experiments.
Total Tokens
We use 2.52.5B tokens for each SAE training.
Appendix H Use ASI on other Activation Functions
H.1 Jumprelu
Another wildly used activation fuction is Jumprelu (Rajamanoharan et al., 2024). We trained the Jumprelu SAEs under two different hyperparameter settings: one with a consistently low $\ell_{0}$ value and another where $\ell_{0}$ gradually decreases from a higher initial value, as described in Appendix G.4. We observed that our method is effective in the former case (Figure 13) but shows little improvement in the latter (Figure 14).
For cases where one follows a schedule that gradually reduces $\ell_{0}$ from a high initial value, we recommend first applying PCA to reduce the dimensionality of the data. The SAE can then be trained on the reduced representation until the $\ell_{0}$ level reaches the target range. Afterwards, the PCA projection matrix can be folded into the model parameters, and training can continue in the original space. This achieves a similar effect without the drawbacks of prolonged training in the high-$\ell_{0}$ regime.
Figure 13:
Figure 14:
H.2 TopK with K Anneal
To enhance the finding in Section H.1, we conduct experiments on a variant of TopK, which sets K to a high value and then lets it decrease during training (He et al., 2024). We find ASI also fails in this case (Figure 15).
Figure 15:
Appendix I Pseudo-code for implementing Active Subspace Init
Below is a PyTorch-style pseudo-code for Active Subspace Initialization.
Use on SAE
# X: activation batch [batch_size, d_model]
# W_E: decoder weight [d_model, d_sae]
# W_D: decoder weight [d_sae, d_model], initialized uniformly
# d_active_subspace: target subspace dimension
# 1. Demean the activations
demeaned_X = X - X.mean(dim=0) # [batch_size, d_model]
# 2. Compute SVD
U, S, V = torch.svd(demeaned_label) # V: [d_model, d_model]
# 3. Take top-d_init singular vectors
proj_weight = V[:, :d_init]  # [d_model, d_init]
# 4. Fold projection into decoder weights
W_D.copy_(W_D[:, :d_init] @ proj_weight.T)
# 5. Init W_E with W_D.T
W_E.copy_(W_D.T)
Use on Lorsa
# X: activation batch [batch_size, seq_len, d_model]
# mhsa: pretrained MHSA module with W_O and W_V
# W_O, W_V: Lorsa decoder/encoder weights
# d_head: dimension of original MHSA heads
# n_heads: number of original MHSA heads
# n_lorsa_heads: number of Lorsa heads
# 1. Run MHSA to capture per-head outputs
Z = mhsa.compute_z(X) # [batch_size, seq_len, n_heads, d_head]
output_per_head = torch.einsum(’bsnh, nhd->bsnd’, Z, mhsa.W_O)
# 2. For each original head
# initialized some lorsa heads into it’s active subspace
rate = n_lorsa_heads // n_heads
for mhsa_index in range(n_heads):
    head_slice = [rate*mhsa_index:rate*(mhsa_index+1)]
    # [B, S, d_model]
    output=output_per_head[:, :, mhsa_index, :]
    output_flat=output.flatten(0, 1) # [B*S,d_model]
    # 3.1 Demean
    demeaned_output=output_flat - output_flat.mean(dim=0)
    # 3.2 Compute SVD
    U, S, V = torch.svd(demeaned_output)
    # 3.3 Take top-d_head singular vectors
    proj_weight = V[:, :d_head]
    # 3.4 Update part of decoder weights W_O
    W_O[head_slice] = W_O[head_slice, :d_head] @ proj_weight.T
    # 3.5 Update part of encoder weights W_V
    head_trans = mhsa.W_V[mhsa_index] @ mhsa.W_O[mhsa_index]
    W_V[head_slice] = W_O[head_slice] @ head_trans.T
The strategy of initialize $W_{V}$ in Lorsa is a method like the tied initialization used in SAEs to ensure alignment between feature encoding and decoding141414”Match” means encoder can be initialized to predict relatively accurate feature activation values for decoder.. This approach has been shown to be crucial for reducing dead features in SAEs (Gao et al., 2024). We think the same thought could also be used to improve the replacement model for MLP (trancoder and cross layer transcoder), which we leave a deeper investigation to future work.